Solve the equation
48. On a train, it took 3/4 hour to travel a distance of 52 1/2 miles. What was the average speed of the train?
Any help with this equation would be Greatly appreciated.

Answer by stanbon(75887)

On a train, it took 3/4 hour to travel a distance of 52 1/2 miles. What was the average speed of the train?
------------------
Formula: distance = (rate)(time)
52 1/2 = rate *(3/4)
rate = 105/2 * 4/3
rate = 35*2 = 70 mph
===============
Cheers,
Stan H.